What approximate volume in mL of 0.200 M NaOH(aq) is needed to form 500 mg of Na2SO4(aq)? The MM of NaOH is 40. g/mol. The MM of Na2SO4 is 142. g/mol.
Answer #1

Required reaction:

2 NaOH + H2SO4 ---------> Na2SO4 + 2H2O

No.of moles of Na2SO4 formed = (mass) / (molar mass) = (0.5 g) / (142 g/mol) = 0.00352 mole

From the balanced reaction, we have :

2 moles of NaOH reacts to form 1 mole of Na2SO4

Hence, 0.00352 mole of Na2SO4 is formed when 2 * 0.00352 mole of NaOH reacted

= 0.00704 mole of NaOH

Therefore volume of NaOH needed is given by:

Volume = (no.of moles) / (molarity) = (0.00704 mole) / (0.200 M)

Volume = 0.03521 L

Volume = 35.21 mL -------- (answer)